# RoPINN: Region Optimized Physics-Informed Neural Networks

### v1

@@ C.1 Benchmarks
Chaotic GS 2D+Time 2 65536 61780 $\boldsymbol{u}_{t}=\varepsilon_{1}\Delta\boldsymbol{u}+b(1-\boldsymbol{u})-\boldsymbol{u}\boldsymbol{v}^{2},\boldsymbol{v}_{t}=\varepsilon_{2}\Delta\boldsymbol{v}-d\boldsymbol{v}+\boldsymbol{uv}^{2}$
@@ C.3 Implementations
As we described in the main text, we set the initial region size $r=10^{-4}$, past iteration number $T_{0}=5$ and only sample 1 point for each region at each iteration for all datasets. The corresponding analyses have been included in Figure 2 for $r$, Figure 3 for sampling points and Appendix D.1 for $T_{0}$ to demonstrate the algorithm property under different hyperparameter settings.
@@ D.1 Hyperparameter Sensitivity on $T_{0}$
As we stated in Algorithm 1, we adopt the gradient variance of past $T_{0}$ iterations to approximate the sampling error defined in Theorem 3.9. In our experiments, we just choose $T_{0}=5$ for all benchmarks, which can already achieve a good performance. To analyze the effect of this hyperparameter, we further add experiments with different choices in Figure 7.
@@ D.2 Experiments with Data Loss (Inverse Problem)
D.3 More Showcases
D.4 Standard Deviations
@@ Appendix G Limitations
This paper presents region optimization as a new PINN training paradigm and provides both theoretical analysis and practical algorithms, supported by extensive experiments. However, there are still several limitations. In the theoretical analysis, we assume that the canonical loss function is $L$-Lipschitz-$\beta$-smooth, which may not be guaranteed in practice. Besides, RoPINN involves several hyperparameters, such as initial region size $r$, and number of past iterations $T_{0}$. Although we have studied the sensitivity w.r.t. them in Figures 2 and Appendix D.1 and demonstrate that they are easy to tune in most cases, we still need to adjust them for better performance in practice. Some hyperparameter tuning tools, such as Weights and Bias (Wandb [^1] https://github.com/wandb/wandb), may mitigate this limitation to some extent, which has already been used in previous related work [34].

### v2

@@ C.1 Benchmarks
Chaotic GS 2D+Time 2 65536 61780 $\boldsymbol{u}_{t}=\varepsilon_{1}\Delta\boldsymbol{u}+b(1-\boldsymbol{u})-\boldsymbol{u}\boldsymbol{v}^{2}$
$\boldsymbol{v}_{t}=\varepsilon_{2}\Delta\boldsymbol{v}-d\boldsymbol{v}+\boldsymbol{uv}^{2}$
@@ C.3 Implementations
As we described in the main text, we set the initial region size $r=10^{-4}$, past iteration number $T_{0}\in\{5,10\}$ and only sample 1 point for each region at each iteration for all datasets. The corresponding analyses have been included in Figure 2 for $r$, Figure 3 for sampling points and Appendix D.1 for $T_{0}$ to demonstrate the algorithm property under different hyperparameter settings.
@@ D.1 Hyperparameter Sensitivity on $T_{0}$
As we stated in Algorithm 1, we adopt the gradient variance of past $T_{0}$ iterations to approximate the sampling error defined in Theorem 3.9. In our experiments, we choose $T_{0}$ from $\{5,10\}$, which can achieve consistently good and stable performance among different benchmarks and PDEs. To analyze the effect of this hyperparameter, we further add experiments with different choices in Figure 8.
@@ D.2 Experiments with Data Loss (Inverse Problem)
D.3 Standard Deviations
@@
D.4 More Showcases
D.5 Experiments with Advanced Quadrature Methods
Table 7: Comparison between Monte Carlo approximation (Monte Carlo) and Gaussian quadrature (Gaussian) on 1D Reaction. rMSE is recorded.
PINN+RoPINN Monte Carlo Gaussian
Sample 1 Point 0.095 0.109
Sample 4 Points 0.066 0.059
Sample 9 Points 0.033 0.030
In our implementation, RoPINN employs a simple Monte Carlo sampling to approximate integral. Obviously, we can adopt more advanced quadrature methods, such as Gaussian quadrature [16]. Thus, we also experiment with 2D space Gaussian quadrature, which requires square number points and the one-point-sampling situation will degenerate to the center value. As shown in Table 7, we can find that under our official setting (only sampling one point), Monte Carlo is better, while Gaussian quadrature is better in more points. Note that although Gaussian quadrature has the potential to achieve better performance, sampling more points may contradict our motivation to boost PINNs without extra backpropagation or gradient calculation. Thus, we choose the Monte Carlo method, which works better under high-efficiency settings.
@@ Appendix F Related Work
Sampling-based methods
Strategies for sampling collocation points perform an important role in training PINN models [46]. Previous sampling-based methods mainly focus on accumulating collocation points to high-residual areas [51, 23] or considering the temporal causality [45], whose theoretical analyses are usually based on the quadrature theorem [32]. Distinct from these methods, RoPINN is motivated by the optimization deficiency of PINN models and can be seamlessly integrated with sampling-based methods with significant promotion (Table 3), indicating that RoPINN works orthogonally to sampling methods. Besides, the theoretical analysis of RoPINN also starts from the optimization perspective, which reveals that one key advancement of RoPINN is a better balance between optimization error and generalization error (Theorem 3.12).
@@ Appendix G Limitations
This paper presents region optimization as a new PINN training paradigm and provides both theoretical analysis and practical algorithms, supported by extensive experiments. However, there are still several limitations. In the theoretical analysis, we assume that the canonical loss function is $L$-Lipschitz-$\beta$-smooth, which may not be guaranteed in practice. Besides, RoPINN involves several hyperparameters, such as initial region size $r$, and number of past iterations $T_{0}$. Although we have studied the sensitivity w.r.t. them in Figures 2 and Appendix D.1 and demonstrate that they are easy to tune in most cases, we still need to adjust them for better performance in practice.
According to our experiments and theorems, we provide some recipes for hyperparameter tuning in the following, which may be helpful to the usage of RoPINN:
  • As shown in Figure 2, region size $r$ will be progressively adjusted by RoPINN. Setting $r$ in $[10^{-6},10^{-4}]$ can work well. According to Theorem 3.12, the choice of $r$ should balance optimization and generalization, which may be inherently decided by the PDE smoothness.
  • As analyzed in Figures 3 and 4, sampling 1-30 points can gain consistent promotion but will linearly increase the computation costs. Following our default setting (sampling one point) can already achieve a competitive performance in a wide range of PDEs.
  • As presented in Figure 8, number of past iterations $T_{0}$ is easy to tune in $[1,20]$. Setting $T_{0}\in\{5,10\}$ can be a good choice, which has been widely verified in our paper.
  • Some hyperparameter tuning tools, such as Weights and Bias (Wandb [^1] https://github.com/wandb/wandb), may mitigate this limitation to some extent, which has already been used in previous related work [37].
Appendix H Broader Impacts
In this paper, we develop a new region optimization training paradigm for PINNs and provide both theorem analyses and practical algorithms. This new perspective may inspire the subsequent research of PINNs, especially rethinking the canonical objective function. In addition, our proposed RoPINN shows favorable efficiency and generalizes well in different base models and PDEs, which can be used to boost the precision of PINNs and generally benefit the downstream tasks, such as physics phenomenon simulation, biological property analysis, etc. Since we purely focus on the training algorithm of PINNs, there are no potential negative social impacts or ethical risks.